OpenStudy (anonymous):

PLEASE HELP! Find all solutions in interval (0,2pi) 2 sin^2 (x) + cos (x) = 1 Write in terms of pi PLEASE HELP!

OpenStudy (anonymous):

PLEASE HELP!!

zepdrix (zepdrix):

\[\Large\rm \sin^2x=1-\cos^2x\]We can rewrite this as a quadratic involving cosx

OpenStudy (anonymous):

use the identities

zepdrix (zepdrix):

\[\Large\rm 2(1-\cos^2x)+\cos x=1\]

zepdrix (zepdrix):

Understand that step? :o

OpenStudy (anonymous):

I would but I don't know what subject this is for

OpenStudy (anonymous):

would the answer be 2pi and 2pi/3. Please someone check it!

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

please someone check it over, this is so important for me

zepdrix (zepdrix):

k sec,

OpenStudy (anonymous):

so sorry but I can't

OpenStudy (anonymous):

thank you zepdrix!!!

zepdrix (zepdrix):

I'm coming up with factors of:\[\Large\rm (\cos x-1)\left(\cos x+\frac{1}{2}\right)=0\] cos x=1 gives us the 2pi, that looks good! cos x = -1/2 should give us a little bit more than just 2pi/3, you're on the right track though.

OpenStudy (anonymous):

so its not 2pi/3 ? :(

zepdrix (zepdrix):